40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2405     Entry time: Sun Dec 13 17:43:10 2009
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: free swinging spectra (vacuum) 

The free swinging spectra of ITMs, ETMs, BS, PRM and SRM were measured under the vacuum-condition. The attachment are measured spectra.

It looks there are nothing wrong because no significant difference appear from the past data and the current data (under atmosperic pressure).

So everything is going well.

Attachment 1: summary_FreeSwinging_vacuum.pdf  1.098 MB  Uploaded Tue Dec 15 00:57:21 2009  | Hide | Hide all
summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf summary_FreeSwinging_vacuum.pdf
ELOG V3.1.3-